OpenStudy (anonymous):

prove the identity: cosx (cot x sin x- tan x sin x)=1 - 2 sin^2x. Should i change it all to cos and sin? or distribute cosx?

OpenStudy (anonymous):

change the cotx and the tanx first. it'll cancel things out and might simplify it for ya

OpenStudy (rulnick):

If you replace cos with A/H, cot with A/O, sin with O/H, etc., it reduces to simple algebra. Canceling produces: cos^2 - sin^2 = 1 - 2 sin^2 cos^2 = 1 - sin^2

OpenStudy (anonymous):

cos(x)((cos(x)(sin(x))/sin(x)) - sin(x)(sin(x)/cos(x)) = cos^2(x) - sin^2(x)

OpenStudy (anonymous):

use the rule sin(x) + cos(x) = 1 to elimante the cos^2(x) and you will find that you have the solution to your problem. I actually miss these problems they were soo easy

OpenStudy (anonymous):

sorry sin^(2)(x) + cos^(2)(x) = 1

OpenStudy (anonymous):

thus cos^(2)(x) = 1 - sin^(2)(x) so we just replace the cos with 1 - sin^(2)(x) and we end up with 1- sin^(2)(x) - sin^2(x) = 1-2sin^(2)(x)

OpenStudy (anonymous):

i dont understand how you are suppose to add everything inside with out common denominators?

OpenStudy (anonymous):

inside the parenthese i mean

OpenStudy (rulnick):

Don't add. The original equation is, by definitions of the trig terms, (A/H) ( (A/O) (O/H) - (O/A) (O/H) ) = 1 - 2 (O/H) (O/H) The left-hand side reduces by distributing the (A/H) and canceling to A^2/H^2 - O^2/H^2 This is cos^2 - sin^2. Then see my solution above.

OpenStudy (anonymous):

you don't add anything inside, you cancel out the fractions by converting cot(x) = cos(x)/sin(x) tan(x) = sin(x)/cos(x) so you have sin(x)cos(x)/sin(x) - sin(x)sin(x)/cos(x) sin(x)cos(x)/sin(x) sin(x)/sin(x) = 1 so you are left with cos(x) - sin(x)sin(x)/cos(x)

OpenStudy (anonymous):

Now distribute the outside term cos(x)(cos(x) - sin(x)sin(x)/cos(x)) = cos(x)cos(x) - sin(x)sin(x)cos(x)/cos(x) cos(x)/cos(x) = 1 thus cos(x)cos(x) - sin(x)sin(x)

OpenStudy (anonymous):

I recommend learning the basic laws of division before tackling these problems

OpenStudy (anonymous):

not to patronize

OpenStudy (anonymous):

After you turn Cotx and Tanx and distribute the cosx we get: \[cosx((cosx/sinx)sinx - (sinx/cosx)sinx) = cosx(cosx - (\sin^2x)/cosx)\] \[= \cos^2x - (Sin^2x)\] remember that \[\cos^2x = 1 - sin^2x\] so we can just plug that in for our \[\cos^2x\] and we get:\[1- \sin^2x - \sin^2x \] combining the two sin^2x wet get \[ 1 - 2\sin^2x\] ta da! now go study :)

OpenStudy (anonymous):

haha oh man i suck at these!

OpenStudy (anonymous):

what are you having problems with specifically? Do you agree that